The textbook for a biology class costs $15 more than twice the cost of a used textbook for college algebra. If the cost of the two books together is $129,find the cost of the biology book.

Answer by rfer(16322)

2x+15+x=129
3x=129-15
3x=114
x=114/3
x=$38 algebra
2x+15=$91 biology